# Efficient Active Imitation Learning with Random Network Distillation

### v1

@@ 4.1 Results
BC 0.883 2455 0.367 - - -
DAgger 0.940 2343 0.450 - - -
Lazy-DAgger 0.939 2314 0.575 20990 312 1238
Ensemble-DAgger 0.952 2489 0.626 18391 13883 2871
RND-DAgger 0.944 2490 0.717 368 708 1214
Table 1: Overall performance: This table illustrates the performance of the different algorithms at the end of the $K$ iterations. It also provides the value of the $nswitch$ variable that accounts for the number of expert interventions.
@@
Table 1 and Figure 6 present the performance metrics across the evaluated environments. To ensure a fair comparison, we report results based on the hyperparameter settings that achieve the highest final performance for each model. Additionally, in the Appendix D, we provide a separate set of results where the hyperparameters are adjusted to align the number of context switches between methods.
@@
Our method, RND-DAgger, demonstrates competitive performance relative to Ensemble-DAgger and Lazy-DAgger, indicating that it effectively learns robust policies across different environments. For example, in the HalfCheetah environment, RND-DAgger achieves a cumulative reward of 2490, compared to 2489 for Ensemble-DAgger and 2314 for Lazy-DAgger. These results highlight that RND-DAgger is capable of discovering efficient policies, matching or surpassing the baselines in terms of overall task success. When analyzing performance relative to the size of the training set, RND-DAgger outperforms the other methods in the early stages of the active imitation learning process. This suggests that RND-DAgger is more effective at focusing on critical states where expert guidance is most needed, thereby gathering valuable feedback more efficiently. By concentrating on key interventions, RND-DAgger accelerates policy improvement and requires fewer training samples to achieve comparable or superior results, making it especially advantageous in scenarios with limited expert availability. In summary, the experimental results demonstrate that RND-DAgger not only achieves strong final performance but also exhibits a more sample-efficient learning curve compared to existing methods, validating its capability to optimize expert interventions and rapidly improve policy quality.
When examining the number of context switches between the different methods, RND-DAgger consistently results in significantly fewer context switches compared to Ensemble-DAgger across all environments. This indicates that RND-DAgger is more stable and requires fewer handovers between the expert and the policy, reducing the burden on the expert. When compared to Lazy-DAgger, RND-DAgger shows mixed results. In the RaceCar environment, RND-DAgger generates fewer context switches, highlighting its efficiency in minimizing expert interventions. However, in the HalfCheetah and Maze environments, RND-DAgger produces a similar number of switches. This similarity in HalfCheetah can be attributed to the use of an oracle expert policy, which performs consistently without generating diverse actions. Under these circumstances, the discrepancy measure used by Lazy-DAgger, which relies on comparing expert and policy actions, remains effective for detecting when to switch. In the Maze environment, despite having a comparable number of context switches to Lazy-DAgger, RND-DAgger achieves a higher final performance. This suggests that even with similar intervention frequencies, RND-DAgger is more effective at leveraging expert feedback to improve the policy. Thus, in scenarios like Maze, where achieving high task performance is critical, RND-DAgger is the better choice due to its ability to make more meaningful corrections and drive the agent towards optimal behavior.
@@
To gain a deeper understanding of when our algorithm requests expert supervision, Figure 7 visualizes the context switches between the current policy and the expert (additional figures are provided in Appendix B). As seen in the figure, both Ensemble-DAgger and Lazy-DAgger query the expert significantly more frequently, leading to a higher number of context switches compared to RND-DAgger. The visualization also reveals that RND-DAgger identifies and focuses on critical areas of the track that are challenging for the agent during early learning stages. Specifically, it requests expert intervention primarily in the problematic zones, such as the bottom section of the track near the bumpers and the obstacle immediately following the speeding ramp (top-left corner of the track). These are regions where the bot initially struggles, making precise interventions crucial for improving performance. This observation indicates that RND-DAgger not only minimizes the number of expert queries but also targets the most relevant segments of the environment where guidance is essential. By focusing on challenging states rather than spreading expert effort across all states, RND-DAgger makes more strategic use of expert supervision, resulting in fewer yet more impactful interventions.
@@
Figure 7: In the RaceCar environment, blue trajectories represent states and actions generated by the agent’s policy, while red trajectories indicate those collected by the oracle. The visualization shows that RND-DAgger predominantly requests expert intervention in challenging sections of the track. This targeted approach results in fewer expert traces compared to baseline methods.
@@
Figure 8: Task performance on RaceCar with a real human as an expert. RND-DAgger achieves similar performance than HG-DAgger with roughly the same amount of demonstrations, but outperforms it in terms of expert time spent.
5 Conclusion
In this work, we presented RND-DAgger, a novel approach to active imitation learning that efficiently minimizes the need for expert interventions by leveraging a state-based measure derived from Random Network Distillation (RND). Unlike traditional methods that rely on action-based discrepancies to detect when to seek expert guidance, RND-DAgger focuses on identifying out-of-distribution states where the agent is most at risk of making errors. This allows our method to selectively request expert feedback only when it is truly necessary, thereby reducing the number of context switches and optimizing the allocation of expert time. Overall, our method provides a step forward in developing more practical and human-efficient imitation learning algorithms, making it a valuable tool for training autonomous agents in complex environments. Future work will explore how to extend RND-DAgger to more challenging tasks and incorporate different forms of expert feedback to further improve learning efficiency and applicability.
@@ 6 Reproducibility Statement
To ensure the reproducibility of our work, we provide detailed pseudo-code in section 2 and section 3. The code, along with necessary dependencies, will be made available to reviewers for evaluation. Upon acceptance, we will release a comprehensive open-source codebase, including all environments, datasets, oracle model checkpoints, active learning algorithms, and a detailed guide on how to reproduce our experiments and results.
@@ Appendix A Experimental Details
    • Threhsold $\lambda$ of OOD detection
@@
    • Threhsold $\tau$ for discrepency measure
    • Threhsold $\chi$ for doubt measure
@@
    • Threhsold $\beta_{H}$ for discrepency measure
    • Threshold $\beta_{R}$ for the backward controlled loop (so the criterion to switch back from expert to autonomous)
@@
With L being a positive factor, a hyperparameter to be tuned. For the second threshold of Lazy-DAgger, we set it to be a factor of the first one: $\beta_{R}=\beta_{H}/L^{*}$, with $L^{*}$ being the hyperparameter tuned. That’s a different method compared to other methods, such as Zhang & Cho (2017) or Hoque et al. (2021) who set their threshold so that approximately 20% of the initial dataset is unsafe, or Ensemble-DAgger Menda et al. (2019) that directly grid search the value. The Table 2(d) summarizes the values used for the grid search.
@@
$\beta_{R}\text{ divisor}$ $[1,1.5,2,3,4]$ 2 1.5 2.5
@@ Appendix B Interactive session visualizations
(a) 0
(b) 1
(c) 2
(d) 3
(e) 4
(f) 5
Figure 10: (a) The agent has control. (b) The OOD measure is triggered (RND), the expert is given control. (c)-(e) The expert demonstrates how to recover. (f) The state is in distribution, the expert keeps the control until the stability window is reached.
@@
On Figure 11, we reported the context switches and full trajectories on both the RaceCar and Maze environments. The results are taken from the same session and the same seed to better compare the figures. From (a) to (b) we can clearly see that Ensemble-DAgger got the most context switches, compared to Lazy-DAgger and RND-DAgger that have a comparable amounts, but we can see that the ones from Lazy-DAgger are more concentrated in specific area whereas is RND-DAgger they seem more spread out. Then, from (d) to (f) we can see the corresponding episodes with expert demonstrations (in red) added to the dataset during that session. We can see that the expert demonstrations are either way longer (the whole episode) or way shorter (only a few frames) for Lazy-DAgger compared to our method, providing more consistency of the demonstrations throughout the map. The difference is clearer on the RaceCar environment. Indeed, both on the context switch figures, we can see that Lazy-DAgger and Ensemble-DAgger are comparable, and that the queries for RND-DAgger are concentrated around only two zones that generated most of the failure cases of the bot.
@@
Figure 11: We conduct a qualitative analysis of the the different query methods. We want to picture if the different measures grasped relevant cases where the learner struggles in our environments. (a)-(d) and (g)-(i): Each dot represents a context switch : a transition from autonomous to expert control, thus the exact point where the learner was said lost. (e)-(h) and (j)-(l): In red, the expert demonstrations, in blue the novice segments.
@@ Appendix C Ablation Study
In this ablation study, we evaluate RND-DAgger, focusing on the additive advantages of three key enhancements: a stability window of size $W$, historical context, and the architecture of the random networks (see Section 2(d) for insights on the values tested). The stability window should increase the quality of the frames by letting the expert finish its demonstration and making sure the learner is fully recovered before letting it the control again. The historical context enriches decision-making by incorporating temporal information from past actions and concatenating them to the current state. This particularly helps in tasks where there is a strong relation between the states. For example, a state where the car is close to a wall is not necessarily a bad state, unless the car is going towards it for several frames in a row. However, a historical context chosen too big is detrimental, because it can overshadow useful dimensions of the sate, and thus preventing the predictor network to seize useful information.
Results are reported in Table 3 and Figure 12. We note that the stability window $W$ helps reducing the context switching, without impacting that much the performance at the end of the sessions of RND-DAgger. In other words, it helps decreasing the expert burden, without impacting the task performance.
@@
Figure 12: Ablation study on our method. RND-DAgger compared to: (light blue) RND without stability window, (orange) bigger historical context, (dark blue) smaller architecture for the random networks.
@@
BC 2455 -
RND-DAgger 2490 708
RND-DAgger - stab. window 2432 1985
RND-DAgger - small 2167 1136
RND-DAgger - historic 2172 554
Lazy-DAgger 2314 312
@@
Stability window focus
@@
On the Figure 13 we report the context switches and session trajectories on the RaceCar environment, at the second DAgger epoch of a given seed with and without a stability window. We can clearly see that the number of context switches increases and a third critical zone appeared at the top right of the map. That is due to the fact that the bot is fairly confident in the top straight line of the course, but the last turn is difficult to manage when the velocity of the car is too high, inducing a query to the expert.There are also more context switches at the entry of that straight line, because the learner has to be precise to pass without any troubles with the walls or the speed ramps on the sides. We conclude that the stability window increases the quality of the demonstrations without impacting the quality of the measure, that still grasps the relevant critical areas of the track.
(a) $W=30$
(b) $W=0$
(c) $W=30$
(d) $W=0$
@@
Figure 13: Ablation study: a focus on the effect of the stability window on the RaceCar environment. We see an increase of the context switches and shorter episode traces as there is no stability window.
Appendix D Additional comparisons
In this section we provide additional comparisons on some experiments obtained during the grid search of our hyperparameters. In Figures 14(a) and 14(b) we aligned the experiments that had comparable context switches throughout the sessions. We see that RND-DAgger clearly outperformed LazyDAgger. On Figures 14(c) and 14(d), we reported the experiments aligned with the best context switches. Again, RND-DAgger outperforms LazyDAgger on the early sessions, while reducing the number of context switches.
(a) Perfs
(b) # context switches
Figure 14: Matched the parameters so that the # context switches are . RND outperforms the rest (higher quality of the frames)
(c) Perfs
(d) # context switches
Figure 15: # context switches is taken to be the best for each method

### v2

@@ 4.1 Results
BC 0.883 $\pm$ 0.029 2455 $\pm$ 424 0.367 $\pm$ 0.043 - - -
DAgger 0.940 $\pm$ 0.026 2343 $\pm$ 209 0.450 $\pm$ 0.074 - - -
Lazy-DAgger 0.939 $\pm$ 0.025 2314 $\pm$ 278 0.575 $\pm$ 0.101 3437 $\pm$ 89 312 $\pm$ 45 1238 $\pm$ 79
Ensemble-DAgger 0.952 $\pm$ 0.018 2489 $\pm$ 108 0.626 $\pm$ 0.045 2785 $\pm$ 41 1452 $\pm$ 134 2871 $\pm$ 94
RND-DAgger 0.944 $\pm$ 0.014 2490 $\pm$ 160 0.717 $\pm$ 0.018 368 $\pm$ 11 708 $\pm$ 130 1214 $\pm$ 25
Table 1: Overall performance: This table illustrates the performance of the different algorithms at the end of the K iterations. It also provides the value of the nswitch variable that accounts for the number of expert interventions. Best, Second best
@@
Table 1 and Figure 6 present the performance metrics across the evaluated environments. To ensure a fair comparison, we report results based on the hyperparameter settings that achieve the highest final performance for each model.
@@
Our method, RND-DAgger, demonstrates competitive performance relative to Ensemble-DAgger and Lazy-DAgger, indicating that it effectively learns robust policies across different environments. For example, in the HalfCheetah environment, RND-DAgger achieves a cumulative reward of 2490, compared to 2489 for Ensemble-DAgger and 2314 for Lazy-DAgger. These results highlight that RND-DAgger is capable of discovering efficient policies, matching or surpassing the baselines in terms of overall task success. When analyzing performance relative to the size of the training set, RND-DAgger outperforms the other methods in the early stages of the active imitation learning process (Fig. 6(b)). This suggests that RND-DAgger is more effective at focusing on critical states where expert guidance is most needed, thereby gathering valuable feedback more efficiently. By concentrating on key interventions, RND-DAgger accelerates policy improvement and requires fewer training samples to achieve comparable or superior results, making it especially advantageous in scenarios with limited expert availability. In summary, the experimental results demonstrate that RND-DAgger not only achieves strong final performance but also exhibits a more sample-efficient learning curve compared to existing methods (Fig. 6(a), 6(b) and 6(c)), validating its capability to optimize expert interventions and rapidly improve policy quality.
When examining the number of context switches between the different methods, RND-DAgger consistently results in significantly fewer context switches compared to Ensemble-DAgger across all environments. This indicates that RND-DAgger is more stable and requires fewer handovers between the expert and the policy, reducing the burden on the expert. When compared to Lazy-DAgger, RND-DAgger shows mixed results. In the RaceCar environment, RND-DAgger generates fewer context switches (Fig. 6(d)), highlighting its efficiency in minimizing expert interventions. However, in the HalfCheetah and Maze environments, RND-DAgger produces a similar number of switches (Fig. 6(e) and 6(f)). This similarity in HalfCheetah can be attributed to the use of an oracle expert policy, which performs consistently without generating diverse actions. Under these circumstances, the discrepancy measure used by Lazy-DAgger, which relies on comparing expert and policy actions, remains effective for detecting when to switch. In the Maze environment, despite having a comparable number of context switches to Lazy-DAgger (Fig. 6(f)), RND-DAgger achieves a higher final performance (Fig. 6(c)).This suggests that RND-DAgger intervention criterion leads to more informative expert interventions and is better able to leverage the expert. Thus, in scenarios like Maze, where achieving high task performance is critical, RND-DAgger is the better choice due to its ability to drive the agent towards optimal behavior.
@@
To gain a deeper understanding of when our algorithm requests expert supervision, Figure 7 visualizes the context switches between the current policy and the expert (additional figures are provided in Appendix B). As seen in the figure, both Ensemble-DAgger and Lazy-DAgger query the expert significantly more frequently, leading to a higher number of context switches compared to RND-DAgger. The visualization also reveals that RND-DAgger identifies and focuses on critical areas of the track that are challenging for the agent during early learning stages. Specifically, it requests expert intervention primarily in the problematic zones, such as the bottom section of the track and the obstacle immediately following the speeding ramp (top-left corner of the track). These are regions where the bot initially struggles, making precise interventions crucial for improving performance. This observation indicates that RND-DAgger not only minimizes the number of expert queries but also targets the most relevant segments of the environment where guidance is essential. By focusing on challenging states rather than spreading expert effort across all states, RND-DAgger makes more strategic use of expert supervision, resulting in fewer yet more impactful interventions.
@@
Figure 7: Visualization of agent’s trajectories (blue) and expert intervention trajectories (red) in RaceCar. RND-DAgger predominantly requests expert intervention in challenging sections of the track, requiring fewer expert interventions overall compared to baseline methods.
@@
Additional experiments Appendix C presents ablation studies of RND-DAgger, including an analysis of the impact of the Minimal Expert Time mechanism (showing how it reduces expert context switches). Appendix D focuses on studying the impact of using Ensemble-DAgger compared to RND-DAgger regarding expert time spent (the metric used in Figure 8(b)), showing how RND-DAgger better minimizes expert burden.
@@
Figure 8: Task performance on RaceCar with a human expert. (a) RND-DAgger achieves similar task performance to HG-DAgger with a comparable number of demonstrations. (b) However, by reducing the need for constant expert monitoring, RND-DAgger significantly lowers the total expert time required.
5 Conclusion and limitations
In this work, we presented RND-DAgger, a novel approach to active imitation learning that efficiently minimizes the need for expert interventions by leveraging a state-based measure derived from Random Network Distillation (RND). Unlike traditional methods that rely on action-based discrepancies to detect when to seek expert guidance, RND-DAgger focuses on identifying out-of-distribution states where the agent is most at risk of making errors. This allows our method to selectively request expert feedback only when it is truly necessary, thereby reducing the number of context switches and optimizing the allocation of expert time. Overall, our method provides a step forward in developing more practical and human-efficient imitation learning algorithms, making it a valuable tool for training autonomous agents in complex environments.
Limitations RND-DAgger requires the expert to immediately take control of the agent, which may be impossible or dangerous in complex real-time environments. Predictive approaches to alert the expert before intervention is an interesting area of improvement. Studying how to scale RND-DAgger to more complex environments such as Partially observable MDPs – e.g. pixel-based environments – would be a natural extension of our work. In such scenarios, noisy-TV problems i.e. environments featuring distracting random or diverse observations, bring important new challenges, which may be addressed through representation learning or explicit high-entropy filtering. Lastly, incorporating richer forms of expert feedback is another valuable future line of work.
6 Acknowledgements
This work was granted access to the HPC resources of IDRIS under the allocation 2024- AD011015218 made by GENCI.
@@ 7 Reproducibility Statement
To ensure the reproducibility of our work, we provide detailed pseudo-code in section 2 and section 3. A comprehensive open-source codebase, including all environments, datasets, oracle model checkpoints, active learning algorithms, and a detailed guide on how to reproduce our experiments and results is available at https://sites.google.com/view/rnd-dagger.
@@ Appendix A Experimental Details
    • Threshold $\lambda$ of OOD detection
@@
    • Threshold $\tau$ for discrepancy measure
    • Threshold $\chi$ for doubt measure
@@
    • Threshold $\beta_{H}$ for discrepancy measure
    • Threshold $\beta_{R}$ for the backward controlled loop (i.e. the criterion to switch back from expert to autonomous)
@@
With L being a positive factor, a hyperparameter to be tuned. For the second threshold of Lazy-DAgger, we set it to be a factor of the first one: $\beta_{R}=\beta_{H}/L^{*}$, with $L^{*}$ being the hyperparameter tuned. That’s a different method compared to other methods, such as Zhang & Cho (2017) or Hoque et al. (2021) who set their threshold so that approximately 20% of the initial dataset is unsafe, or Ensemble-DAgger Menda et al. (2019) that directly grid searched the value. The Table 2(d) summarizes the values used for our grid search.
@@
$\beta_{R}\text{ divider}$ $[1,1.5,2,3,4]$ 2 1.5 2.5
@@ Appendix B Interactive session visualizations
On Figure 10, we reported the context switches and full trajectories on both the RaceCar and Maze environments. The results are taken from the same session and the same seed to better compare the figures. From (a) to (b), it is clear that Ensemble-DAgger exhibits the most context switches compared to Lazy-DAgger and RND-DAgger, which show comparable counts. However, the context switches in Lazy-DAgger are more concentrated in specific areas, whereas those in RND-DAgger are more evenly distributed. Then, from (d) to (f) we can see the corresponding episodes with expert demonstrations (in red) added to the dataset during that session. We can see that the expert demonstrations are either way longer (the whole episode) or way shorter (only a few frames) for Lazy-DAgger compared to our method, providing more consistency throughout the map. The difference is clearer on the RaceCar environment. Indeed, on both context switch figures, we see that Lazy-DAgger and Ensemble-DAgger are comparable, and that the queries for RND-DAgger are concentrated around only two zones that generated most of the failure cases.
@@
Figure 10: We conduct a qualitative analysis of the the different query methods. (a)-(d) and (g)-(i): Each dot represents a context switch : a transition from autonomous to expert control. (e)-(h) and (j)-(l): In red, the expert demonstrations, in blue the novice segments.
(a) 0
(b) 1
(c) 2
(d) 3
(e) 4
(f) 5
Figure 11: (a) The agent has control. (b) The OOD measure is triggered (RND), the expert is given control. (c)-(e) The expert demonstrates how to recover. (f) The state is in distribution, the expert keeps the control until the Minimal Expert Time is reached.
@@ Appendix C Ablation Study
C.1 RND-DAgger on HalfCheetah
In this ablation study, we evaluate RND-DAgger, focusing on the additive advantages of three key enhancements: a Minimal Expert Time (MET) of size $W$, historical context, and the architecture of the random networks (see Section 2(d) for insights on the values tested). The addition of a Minimal Expert Time should increase the quality of the frames by letting the expert finish its demonstration and making sure the learner is fully recovered before letting it the control again. The historical context enriches decision-making by incorporating temporal information from past actions and concatenating them to the current state. This particularly helps in tasks where there is a strong relation between the states. For example, a state where the car is close to a wall is not necessarily a bad state, unless the car is going towards it for several frames in a row. However, a historical context chosen too big is detrimental, because it can overshadow useful dimensions of the sate, and thus preventing the predictor network to seize useful information.
Results of that ablation are reported in Table 3 and Figure 12. We note that the Minimal Expert Time $W$ helps reducing the context switching, without impacting that much the performance at the end of the sessions of RND-DAgger. In other words, it helps decreasing the expert burden, without impacting the task performance.
@@ C.1 RND-DAgger on HalfCheetah
Figure 12: Ablation study on our method. RND-DAgger compared to: (light blue) RND w/o MET, (orange) bigger historical context, (dark blue) smaller architecture for the random networks.
@@
BC 2455 $\pm$ 424 -
RND-DAgger 2490 $\pm$ 160 708 $\pm$ 130
RND-DAgger - MET 2432 $\pm$ 175 1985 $\pm$ 215
RND-DAgger - small 2167 $\pm$ 140 1136 $\pm$ 76
RND-DAgger - historic 2172 $\pm$ 290 554 $\pm$ 90
Lazy-DAgger 2314 $\pm$ 278 312 $\pm$ 45
@@
C.2 Minimal Expert Time focus
Ablation study on all the methods
We further investigated the impact of the Minimal Expert Time (MET) on our method and on our baselines, across all the tasks. We reported the results of that study in Figure 13 and Table 4. In particular, we can see that LazyDAgger + MET isn’t appropriate in HalfCheetah. The combination of the Lazy mechanism (introduction of a second threshold to create a hysteresis band) and the MET forces the expert to always have control over the agent, which explains the extremely low value for the context switches (which happens only once per episode) and the poor performance learning curve (Fig. 13(b)). The same remark holds for Ensemble-DAgger on RaceCar, but in this case, this behavior doesn’t seem to impact the task performance of the method (Fig. 13(a)).
(a) Task performance RaceCar
(b) Task performance HalfCheetah
(c) Task performance Maze
(d) Context switches RaceCar
(e) Context switches HalfCheetah
(f) Context switches Maze
Figure 13: Results for the ablation study with focus on the addition of the Minimal Expert Time (MET) on the baselines Ensemble-DAgger and LazyDAgger. RND-DAgger remains the same as in Table 1
Task Performance # Context Switch
RC HC Maze RC HC Maze
Lazy-DAgger + MET 0.943 $\pm$ 0.011 2374 $\pm$ 325 0.575 $\pm$ 0.101 759 $\pm$ 39 78 $\pm$ 0.0 876 $\pm$ 71
Lazy-DAgger 0.939 $\pm$ 0.025 2314 $\pm$ 278 0.590 $\pm$ 0.057 3437 $\pm$ 89 312 $\pm$ 45 1238 $\pm$ 79
Ensemble-DAgger + MET 0.935 $\pm$ 0.019 2502 $\pm$ 228 0.707 $\pm$ 0.023 77 $\pm$ 4 315 $\pm$ 84 2034 $\pm$ 79
Ensemble-DAgger 0.952 $\pm$ 0.018 2489 $\pm$ 108 0.626 $\pm$ 0.045 2785 $\pm$ 41 1452 $\pm$ 134 2871 $\pm$ 94
RND-DAgger 0.944 $\pm$ 0.014 2490 $\pm$ 160 0.717 $\pm$ 0.018 368 $\pm$ 11 708 $\pm$ 130 1214 $\pm$ 25
RND-DAgger w/o MET - 2432 $\pm$ 175 - - 1985 $\pm$ 215 -
Table 4: We conduct an ablation study on the impact of the Minimal Expert Time (MET) on all baselines across our three environments. We report the final task performance and number of context switches. We conducted the same grid search for the value of W as for RND-DAgger (see section A for more details).
Qualitative interpretation
@@ C.2 Minimal Expert Time focus
On Figure 14 we report the context switches on the RaceCar environment, at the second DAgger epoch of a given seed with and without a MET. We can clearly see that the number of context switches increases and a third critical zone appeared at the top right of the map for RND-DAgger (Fig. 14(a)). We interpret that the bot is fairly confident in the top straight line of the course, but the last turn is difficult to manage when the velocity of the car is too high, inducing a query to the expert. There are also more context switches at the entrance of that straight line, because the learner has to be precise to pass without touching the walls or the speed ramps on the sides. We conclude that the MET increases the quality of the demonstrations without impacting the quality of the measure, that still grasps the relevant critical areas of the track.
(a) RND-DAgger $W=0$
(b) Ensemble-DAgger $W=0$
(c) LazyDAgger $W=0$
(d) RND-DAgger $W=30$
(e) Ensemble-DAgger $W=30$
(f) LazyDAgger $W=10$
@@
Figure 14: Ablation study: a focus on the effect of the MET on the RaceCar environment. We see an increase of the context switches and shorter episode traces as there is no Minimal Expert Time for all the methods. RND-DAgger seems to have a better grasp on the difficult zones of the track, whereas the two other measures trigger the expert rather uniformly across the track.
Appendix D Total expert time comparisons
While focusing solely on task performance and context switches provides important insights on the efficiency of a method and on a given task, it doesn’t tell the complete story about expert burden. We conducted an additional analysis on the Total Expert Time, as we did in the Fig 8(b) on human experiments. The Table 5 shows that Ensemble-Dagger is highly time consuming for the expert without statistically outperforming RND-Dagger. For example, on Maze, it is three times longer to achieve a similar performance than with RND-DAgger, even with the addition of our MET mechanism on that baseline. This demonstrates that we are more efficient in real human expert scenarios. Moreover, and especially in a fast-paced environment like RaceCar, that baseline - and its improved MET variant - requires the expert to input actions at the same time as the agent to be able to compute the discrepancy measure, which is not natural in practice.
Task Performance Total Expert Time
RC HC Maze RC HC Maze
RND-DAgger 0.944 $\pm$ 0.014 2490 $\pm$ 160 0.717 $\pm$ 0.018 16664 $\pm$ 73 (27.8 $\pm$ 0.1 mins) 11968 $\pm$ 107 (6.60 $\pm$ 0.06 mins) 16173 $\pm$ 44 (27.0 $\pm$ 0.1 mins)
Ensemble-DAgger 0.952 $\pm$ 0.018 2489 $\pm$ 108 0.626 $\pm$ 0.045 31394 $\pm$ 506 (52.3 $\pm$ 0.8 mins) 21535 $\pm$ 774 (11.8 $\pm$ 0.4 mins) 51434 $\pm$ 1507 (85.7 $\pm$ 2.5 mins)
Ensemble-DAgger + MET 0.935 $\pm$ 0.019 2502 $\pm$ 228 0.707 $\pm$ 0.023 18216 $\pm$ 165 (30.4 $\pm$ 0.3 mins) 13474 $\pm$ 631 (7.40 $\pm$ 0.35 mins) 48125 $\pm$ 900 (80.2 $\pm$ 1.5 mins)
Table 5: The results considering the Total Expert Time i.e. the total number of frames on which the expert had to be involved to provide an action. In the case of Ensemble-DAgger, the expert action is needed at each time step to compute the discrepancy measure between the mean action of the ensemble of policies and the expert action. The time spent in minutes is computed with a typical frame rate of 30 for HalfCheetah (arbitrary) and 60 with one action taken every 6 frames, for the two game environments.